OpenStudy (anonymous):

What is the difference of four and five over eight minus three and one over four in simplest form?

OpenStudy (anonymous):

a. 2 b. 1 c. one and three over eight d. one and three over four

OpenStudy (anonymous):

i suck at fractions I:

OpenStudy (anonymous):

\[4\frac{ 5 }{ 8 } - 3\frac{ 1 }{ 4 } -> simplify.\]

OpenStudy (snowcrystal):

do u know how to do common denomanator

OpenStudy (anonymous):

\[4\frac{ 5 }{ 8 }-3\frac{ 2 }{ 8 }\]

OpenStudy (anonymous):

so c?

OpenStudy (snowcrystal):

to make it easier to subtract this is what you will want to do \[4\frac{ 5 }{ 8 } - 3\frac{2}{8}\] 4 x 8= 32+ 5= 37 your 1st fraction would be \[\frac{ 37 }{ 8 }\] 3 x 8=24+ 2 = 26 your second fraction would be \[\frac{ 26 }{ 8 }\]

OpenStudy (snowcrystal):

what is 37-26?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

You're right, it is C.

OpenStudy (snowcrystal):

correct now the denominator would stay the same your new fraction would be \[\frac{ 11 }{ 8 }\]

OpenStudy (anonymous):

\[\frac{ 11 }{ 8 } -> 1\frac{ 3 }{ 8 }\]

OpenStudy (snowcrystal):

@fathomboy remember we cant just give them the answer we have to explain it so that they can understand

OpenStudy (anonymous):

a. 2 b. 1 c. one and three over eight d. one and three over four